Minutes of the 7th Committee Meeting of the Dublin University Rifle Club

Moderator Minutes taker Mark Dennehy Present Benjamin Jones (Treasurer),

Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer), Conor McDermottroe (Public Relations Officer),

Absent(with apologies) Steven Watterson (Captain),

Absent Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 4 September 2001 1830 – 2020 Distribution DURC Committee

Overview of topics

1 Secretarys Report 1

2 A&E Report 2

3 Rest of Meeting 2 3.1 East Antrim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2 New Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.3 Junior Match in Bisley . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1 Secretarys Report

Mark reported that the TTT was underway with a slow start. Stickers had to be resent to RRPC owing to a mixup in postal addresses, the RRPC contact having moved house and dropped out of RRPC since the last TTT competition.

2 A&E Report

Task: Proceed with the TTT (Mark)  1850  John

Keeney arrives

The Freedom of Information Act applies to college activities as of this October: Mark brought the advisory notes from College and they were discussed at length. It was agreed that the Club would have to pay more attention to its records. It was also agreed to add a warning to email regarding non-confidentiality of email. Task: Append warning to email signatures regarding non-confidentiality (All)

2 A&E Report

Pat Moray is still not happy with the situation as regards the storage of privately held rifles in college facilities.

Millards no longer have any Lapua Master ammunition in stock. We need to select ammunition to use for the coming year. Winchester T22 remains the standard training ammunition grade, but competition grade has to be selected as Lapua are approaching their 60% price rise. It was agreed to get a price for the new SK-Lapua ammunition. Task: Price SK ammunition (John)

3 Rest of Meeting

At this point the meeting structure broke down and so the rest of the minutes are listed as the discussion progressed

Ben reported that Cathy is unhappy about the level of transparancy in the procedure for purchasing ammunition and is concerned over DUCAC paying for some ammunition twice. This poses the problem of raising sufficent cash to pay for ammunition on the order of 10,000 rounds which is the quantity Millards wish to deal in. Cathy says DUCAC will pay for competition ammunition but she wants paperwork for each competition detailing who has shot how much ammunition.

This was agreed to be a major problem - the amount of money required would cripple the Club as we do have to purchase consumables to operate. It was agreed that social events have to be cut back for financial reasons and that corporate sponsorship needed to be sought as previously discussed.

Some aspects of Freshers Week were briefly discussed, it being noted that the traditional party needed to be cut back, the safety lecture shortened and the possibility of upping the membership fees was mentioned but dismissed.

Ben pointed out that this is the start of the College Financial Year and thus our grants are coming up. It was agreed to defer that discussion however.

Mark suggested that if we are to be audited, that the club pursue the payment of the remainder of the monies owed by DUCAC over the purchase of the Kurt Thune jackets, which was several hundred pounds. John was convinced that that money was not retrieveable, Mark was convinced it was if the club did not pursue it. The matter was not resolved.

3 Rest of Meeting

Ben wanted an offical decision on whether we were going to buy bulk ammunition - it was decided that we would not do so yet.

Ben noted the high level of RO Credits, a system currently being abused by ROs, and asked the highest debtors to pay promptly. Alan said he would pay in full on Sept. 20th, his first payday in his new Postgraduate post, and Conor paid £15 on the spot to clear his account. Task: Pay RO Credits promptly (All)

Mark reported that there was a possibility that we would receive a bill from FSC for range hire for the DURC Open. This brought up the matter of bills for last years events. A letter had been sent offically to the NTSA regarding the paperwork required for billing DURC, and imposing a deadline of August 31st for receipt of bills for last years events. If we were to pay the FSC bill, it was felt this would invalidate that notice. A proposal was put forward that all bills received after August 31 for events before August 31 would be summarily refused. An offical vote was held : Mark, Conor and Ben voted for, Alan and John were opposed. The proposal was therefore carried. 1930 John left

3.1 East Antrim

The paperwork was reported as being complete for going to the EARC Festival on September 29/30. Mark’s paperwork for his personal rifles is not complete owing to Civil Service delays. Transport arrangements were discussed - Ray, Peter and Alan are able to drive in N.Ireland. Mark cannot due to insurance restrictions by AXA and he cannot have anyone other than Ray or Keith on his insurance due to an age restiction by AXA. It was hinted that he might want to try for his full licence - it was then pointed out by Mark that that was a strange request since all present had great expertise in pointing out flaws in his driving technique ☺. The list of people going was Mark, Ray, Ben, Alan, Peter, Steven and John, with Liam Murray as a possible. Conor cannot go barring him finding someone to cover at work for that weekend - he was encouraged to do so due to the size of the EARC competition.

3.2 New Rules

Mark brought up two new rules from the NTSA meeting. Firstly was a safety ruling regarding the firing of shots into the backstop at Rathdrum - owing to the nature of that backstop, this was a dangerous procedure and it will be penalised in future.

Second was a ruling on claiming cards at competitions - it is now offically ruled that if you claim your cards at a competition, you formally give up your right to challenge scores and you lose all splits by default. This is due to the nature of the regauging process used in determining splits. Procedures for shooting competitions were discussed in light of this ruling, though Mark pointed out that this was standard at all NSRA competitions.

3.3 Junior Match in Bisley

The fiasco of organisation that this match turned out to be was discussed - Mark had made an offical complaint to the NTSA and had been given apologies by both Pat O’Brien and Geoff Cooney. Ben and Alan wanted a formal complaint made, however Mark pointed out that this would be a poor mark on Geoffs record when he is about to bring in a large and much-needed change to the classification system. It was agreed that Mark would as ask for a written apology instead. Task: Ask Geoff and Pat for a written apology for DUCAC’s eyes (Mark)

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

Tasks agreed on at meeting

Mark Proceed with the TTT (4 September 2001) [/133] . . . . . . . . . . /2 All Append warning to email signatures regarding non-confidentiality

(4 September 2001) [/147] . . . . . . . . . . . . . . . . . . . . . . . . . /2

John Price SK ammunition (4 September 2001) [/160] . . . . . . . . . . . /2 All Pay RO Credits promptly (4 September 2001) [/209] . . . . . . . . . /2 Mark Ask Geoff and Pat for a written apology for DUCAC’s eyes (4

September 2001) [/275] . . . . . . . . . . . . . . . . . . . . . . . . . . . /2